OpenStudy (anonymous):

NEED HELP: ( I will solve, just need help) If h(x) = x - 7 and g(x) = x², which expression is equivalent to (g o h)(5)? (5 - 7)² (5)² - 7 (5)²(5 - 7) (5 - 7)x² I know that (g o h)(5) = g(h(5)) and I need to find h(5) That would be h(5) = 5 - 7 But after that I am lost. *The answer I had down was (5 - 7)x² because I didn't think you could do anything else after the (5 - 7)

OpenStudy (cggurumanjunath):

h(5)=5-7=-2

OpenStudy (cggurumanjunath):

g(h(5))=g(-2)

OpenStudy (cggurumanjunath):

g(-2)=(-2)^2=4

OpenStudy (anonymous):

But in all of the answers, I don't see where you could get 4 from. I understand that h(5) = 5-7 = -2 and g(-2) = -2² = 4 but I don't see how I would find that in the answers.

OpenStudy (amistre64):

gox, or rather, g "composed of" x, is equal to: x^2 goh, or rather, g "composed of" h, is equal to: h^2 but h "composed of" x is equal to: x-7 so g(h(x)) = (x-7)^2

OpenStudy (anonymous):

Ok I gotcha. Thank you!